Soft drinks are commonly sold in aluminum containers. How many such containers are thrown away or recycled each year by consumers ( Assume an average 3 cans per person each week and a population of 320 million. Enter your answer using one of the following formats 1.2e-3 for 0.0012 and 1.20e+2 for 120. )


1
Expert's answer
2021-02-03T16:12:09-0500

There are approximately 51 weeks in year. Thus, each person throw away "51\cdot 3 = 153" cans per year. The population of 320 million uses


"153\cdot 320\times 10^6 \approx 4.9\times10^{10}"

cans per year.


Answer. 4.9e+10.